OpenStudy (anonymous):

8. What is the total measure of all of the interior angles in a convex 15-gon?

OpenStudy (anonymous):

i've forgotten the formula but you can work it out this way average measure of external angle = 360 / 15 = 24 i internal = 180-24 = 156 so total = 156 * 15 = 2340 degrees

OpenStudy (anonymous):

soo 156?

OpenStudy (anonymous):

total measure of all the angles = 2340

OpenStudy (anonymous):

156 is the measure of one angle in a regular 15-gon

OpenStudy (anonymous):

by the way re-the answer for exterior angle in 18-gon that emerald queried - in my day it was 20 degrees -his definition gives 200 degrees. Maybe you had better check it out with your math instructor.